OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (anonymous):

would tho be -10, -8, -6,... ETC?

OpenStudy (jdoe0001):

-10 + (2) = -8 -10 +( -2) = -12

OpenStudy (anonymous):

oh yea my bad

OpenStudy (anonymous):

I dont know why but when it comes to negative numbers for some reason i always get the problem wrong :/ But hey thanks!

OpenStudy (jdoe0001):

yw